A spherically symmetric planet has 2 times the Ear
Answer #1

Gravitational force is given by :

F = G Mm/r^2 , M = earth mass, m = jar mass, r = distance betweeen them

given for earth F = 12N

planet mass = 2M , planet radius = 2r

F_planet = G*2M*m/(2r)^2   = GMm*2/4r^2 = 0.5 GM m/r^2 = 0.5 F

hence the force on planet is half that on earth, so weight will also be half , ie = 6 N